### v1

# Ab Initio Framework of Electron-Phonon Coupling for Chiral Phonons with Giant Phonon Magnetic Moments in Magnetic Materials

Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, this field remains severely limited due to the inability to accurately account for electron-phonon-coupling (EPC) effects for chiral phonons. Here, we resolve this challenge through developing an ab initio framework for computing EPC-driven phonon magnetic properties, applicable to all insulating and metallic materials. Our theory predicts a dramatic enhancement of phonon magnetic moments exceeding classical predictions by multiple orders, accompanied by distinct Zeeman splitting. Interestingly, these splittings induce observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states. Through first-principles calculations, we identify five topological phonon magnets with intrinsic phonon Chern states and also propose to detect low-energy neutral particles such as dark matters. This research not only establishes a computational framework for investigating EPC-driven phonon magnetic properties, but also provides a material design strategy for realizing topological phonon states in quantum magnetic materials.
Introduction. Chiral phonons, defined as collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and his cooperators [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. These chiral vibrations not only reveal fundamental properties of phonons previously unrecognized, but also establish a novel degree of freedom for manipulating phonon-mediated phenomena [1, 4, 5, 6, 7]. Recently, there has been a surge of research interest in chiral phonons, spanning theoretical investigations [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] to experimental explorations [19, 20, 21, 22, 23, 24, 25, 26, 27]. Of particular importance are phonon magnetic moments (PMMs) and their associated phonon Zeeman splittings, which constitute key mechanisms in chiral-phonon-involved effects [6, 7]. However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below practical application requirements [13], posing significant limitations for the advancing development and application of chiral phonons. Although electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [11, 13, 16], a unified ab initio theoretical framework of EPC-driven PMMs for realistic materials is lacking.
Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-driven PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical Zeeman splitting of chiral phonons is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
Besides magnetic moments, phonons can also demonstrate the Hall effect [28, 29, 30, 31] and manifest topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34]. When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, essentially realizing phonon Chern insulators. These topological phonon states featuring topologically protected edge currents could significantly enhance both detection sensitivity and practical exploitation of the phonon Hall effect. Nevertheless, conventional TRS-breaking approaches for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations or induce parasitic effects [35, 36, 37]. Consequently, phonon Chern states driven by intrinsic magnetic moments without external fields are highly desirable. Crucially, predicting such intrinsic topological phonon states in magnets demands precise calculations of TRS-broken phonon spectra, which remains challenging due to the lack of ab initio framework for magnetic phonon systems [38, 39, 40, 41, 42, 43].
To resolve these challenges, we establish a unified ab initio framework for determining EPC-driven phonon magnetic properties, which enables first-principles calculations of the previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Our theory demonstrates EPC-driven enhancements of PMMs by four orders of magnitude, inducing giant chiral phonon Zeeman splittings. Building on these giant splittings, we propose the inertially decoupled model to realize intrinsic phonon Chern states in magnets. Furthermore, through first-principles calculations, we identify five promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$. Our work not only paves a material-design pathway for topological phonon magnets, but also establishes the framework for chiral phonon physics in magnetic phenomena.
Ab initio framework for phonon Zeeman energy. Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [11, 13, 14, 16]. However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop a unified theory that can comprehensively describe phonon behaviors in insulators, semimetals, and metals.
We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [44], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$. Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass. By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$. Here, we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$. Therefore, the Zeeman energy of the chiral phonon can be regarded as an effective linear response to its angular momentum.
We notice that $\bm{\eta}$ can be deduced from the TRS-breaking self-energy correction of chiral phonons driven by EPC. Remarkably, within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b)(see the supplement material (SM)) [45],
$\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (1)
Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively. $N$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$. $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [46, 47]. Notably, all parameters in Eq. (1) can be directly derived from first-principles calculations, obviating the need for empirical inputs or additional assumptions. As a result, the previously inaccessible TRS-breaking phonon spectra driven by EPC in both insulating and metallic systems is now accessible across the whole Brillouin zone.
Phonon Zeeman splitting and enhanced phonon $g$-factor. We propose the essential conditions for realizing significant phonon Zeeman splittings by EPC. Since the primary determinant of the phonon Zeeman splitting is the susceptibility $\bm{\chi}^{-}$, there are two key points according to Eq. (1): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$. (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, the parametric sensitivity in Eq. (1) leads to an enhancement of $\bm{\chi}^{-}$ [45]. Consequently, we can conclude that EPC-driven large phonon Zeeman splittings are most likely to occur in narrow-gap semiconductors, semimetals and metals comprising light elements.
Large phonon Zeeman splittings suggest that phonons exhibit significant phonon magnetic moments. To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [45]),
$\mu^{\text{ph}}_{c}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (2)
Here, $\mu_{B}$ denotes the Bohr magneton, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon, where $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$, $g^{e}_{c}$ is the $g$-factor of electrons along $c$ direction involved in EPC, and the dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor. This amplification factor is primarily influenced by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electronic states, and it can significantly exceed unity. This enhancement allows the $g$-factor of the chiral phonon to the order of $g^{e}_{c}$.
This EPC-enhanced PMM can be intuitively understood through the following mechanism. As chiral phonons drive atomic cyclotron motion, a portion of its angular momentum is transferred to the electrons via EPC. Owing to the considerably small mass of electrons, the angular momentum transfer gives rise to strong loop electron currents, thereby resulting in a significant PMM. In contrast, within the classical framework, the phonon’s $g$-factor is constrained to the order of $10^{-5}\sim 10^{-4}g^{e}_{c}$ due to heavy ionic masses compared to the electron mass [13, 48].
Figure 2: Inertially decoupled model (IDM) and intrinsic phonon Chern states. (a) Schematic of optical chiral phonon modes: When $m\sim M$, both atoms contribute to the high-frequency model (up); in the IDM condition ($m\ll M$), the optical model is primarily contributed by light atoms (down). (b) Phonon dispersions of nonmagnetic honeycomb lattice and magnetic triangle lattice. (c-e) TRS-breaking phonon dispersions of 2D/3D IDM structures: the 2D IDM by combining the honeycomb and triangular lattices (c,d), and the 3D IDM by stacking 2D IDM along the $z$-direction (e). The color gradient from violet and yellow represents the contributions from the honeycomb and triangular lattice, and the colored gradient from blue and red represents the phonon angular momentum magnitude. (f) Edge states (cyan/magenta) of IDM under open boundaries. Parameters are detailed in the SM [45].
Intrinsic phonon Chern states and the inertially decoupled model. The large phonon Zeeman splitting can pave the way toward a phonon Chern state through the phonon band inversion [45]. Following the analogy to the Chern state in the Haldane model [49], we adopt the two-dimensional (2D) honeycomb lattice [44, 50, 51] as an illustrative example. When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $m_{I}$ and $m_{T}$, respectively. It has been shown that when $|m_{T}|>|m_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [44].
However, realizing such phonon Chern states in realistic materials hosting honeycomb lattices remains challenging. Although Coriolis forces and strong magnetic fields can produce a nonzero $m_{T}$ [44], these approaches either require large external fields or complex experimental measurements. In contrast, the EPC-enhanced phonon Zeeman splitting could offer a sizable $m_{T}$ in magnets without the need of external fields. To further satisfy the condition $|m_{T}|>|m_{I}|$, it is necessary to minimize $|m_{I}|$. Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|m_{I}|=0$ would be the ideal candidate. Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered. Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $m_{I}$, which will make the condition $|m_{T}|>|m_{I}|$ difficult to satisfy.
To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference. A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs. 2(b,c). Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions mainly originate from the light and heavy atoms, respectively, as shown in Figs. 2(a,c), which is termed “inertially decoupled” (see SM [45] for details). Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $m_{I}$ mass at both the $K$ and $K^{\prime}$ valleys. Meanwhile, the magnetic $M$ atoms breaks TRS and provide effective magnetic fields to generate large Zeeman splittings of chiral phonons via EPC, producing the desired large $m_{T}$ mass. The large $m_{T}$ opens a topologically nontrivial phonon gap (seen in Fig. 2(c,d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by the topological edge states crossing the phonon gap, as shown in Fig. 2(f).
The inertially decoupled model is also applicable to three-dimensional (3D) magnets. Interestingly, when the nonmagnetic honeycomb lattice (the $m$ sublayer) and the magnetic triangular lattice (the $M$ sublayer) are stacked alternately along the $z$-direction, each $M$ sublayer effectively acts as a buffer layer that nearly decouples the phonon dispersions of adjacent $m$ sublayers in the high-frequency regime. A representative phonon dispersion of such a 3D bulk model is shown in Fig. 2(e). Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits an intrinsic 3D phonon Chern state. This configuration hosts multiple chiral phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$. The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$. (c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the TRS breaking, respectively. (e) The phonon dispersion of $\mathrm{EuSi_{2}}$ under open-boundary conditions. (f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the TRS breaking, respectively. (h) The phonon dispersion of the sandwiched slab under open-boundary conditions. The color gradient from red to blue represents the magnitude of the phonon angular momentum. Cyan and magenta label the surface or edge modes originating from opposite boundaries, respectively.
Candidate materials. Building upon the above theoretical framework, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [46, 47] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [45]), yielding TRS-broken phonon dispersions.
Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal $P6/mmm$ magnets, specifically $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [58], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [59], and $\mathrm{CrB_{2}}$ [60, 61], exhibiting intrinsic phonon Chern states via EPC-enhanced Zeeman splittings ($\Delta_{\text{z}}=2|E_{\text{Z}}|$). As summarized in Table I, they all have large observable phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\gamma$ is phonon linewidth resulting from the EPC [62, 63]. Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
Here, we take $\mathrm{EuSi_{2}}$ as a representative example (see SM [45] for details on all candidate materials). The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs. 3(a, e). Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a magnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction. The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$. As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs. 3(c, d)], opening a sizable full gap ($\approx 0.35$ THz) throughout the Brillouin zone. Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state. This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig. 3(e). To achieve a 2D Chern state, we construct a sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig. 3(b). By comparing Figs. 2(f) and 2(g), we can see that the Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ in the phonon dispersion of this sandwiched slab, and the accompanying single pair of edge modes traversing the phonon gap, as shown in Fig. 3(h), provide definitive evidence for intrinsic phonon Chern states.
Table 1: The direct phonon gaps at K points for the five candidate materials, where $\gamma$ and $\Delta_{\text{z}}$ represent the phonon linewidth and the phonon Zeeman splitting, respectively. The observable phonon gap $\Delta_{\text{ph}}$ is defined as $\Delta_{\text{z}}-\gamma$.
Mat. $\Delta_{\text{z}}$ (THz) $\gamma$ (THz) $\Delta_{\text{ph}}$ (THz)
$\mathrm{EuSi_{2}}$ 0.426 0.049 0.377
$\mathrm{GdSi_{2}}$ 0.406 0.061 0.345
$\mathrm{EuGa_{2}}$ 0.134 0.028 0.106
$\mathrm{GdGa_{2}}$ 0.145 0.015 0.130
$\mathrm{CrB_{2}}$ 0.270 0.033 0.237
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring extreme external strong magnetic fields, Coriolis forces, or stress gradients. One potential advantage may be its ability to detect neutral particles by converting them into phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within that of the first “conduction” band above the topological gap in the phonon dispersion, this process may potentially excite the corresponding phonon modes [64, 65], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via edge states, producing a edge phonon current immune to scattering[Fig. 4(a)]. This robust edge current is significantly more measurable than conventional phonon excitations dispersed throughout the material. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, the gap determines the maximum operating temperature for such a detector. Fig. 4(b) lists the transfered energy ranges and gaps for all five candidate materials (see SM [45] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to light-mass dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates.
Figure 4: Detectors of neutral particles based on the intrinsic phonon Chern state. (a) Schematic illustrations of the detection process. Neutral particles excite phonon modes in high-frequency “conduction” bands. These modes then gradually relax to lower-frequency regimes via chiral edge phonon channels. (b) Transfered energy ranges and phonon Chern gaps for all the layered candidate materials.
In summary, we have established a unified ab initio framework enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems. Under this framework, we propose an inertially decoupled model to design intrinsic phonon Chern states, and identify five candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through phonon edge current transduction. These advances simultaneously open broad pathways for revealing new phonon magnetic effects in quantum magnetic materials.
Acknowledgements.
The authors thank Lei Zhang and Zuowei Liu for valuable discussions. This work is partly supported by National Key Projects for Research and Development of China (Grants No. 2024YFA1409100, No. 2021YFA1400400 and No. 2022YFA1403602), the Natural Science Foundation of Jiangsu Province (No. BK20220032, BK20233001, BK20243011), the Natural Science Foundation of China (No. 12074181, No. 12104217, No. 12174182, and No. 92365203), and the e-Science Center of Collaborative Innovation Center of Advanced Microstructures.